# Vessel Detection and Localization Using Distributed Acoustic Sensing in Submarine Optical Fiber Cables

### v1

@@ Abstract
Index Terms:
Vessel detection, vessel localization, Distributed Acoustic Sensing, Submarine Cables.
@@ I Introduction
Submarine cables, essential for global internet connectivity, energy transmission, and communication, are particularly vulnerable to accidental damage (e.g., fishing nets, ship anchors) and sabotage linked to geopolitical tensions [1]. Recent incidents highlight this risk: in November 2024, damage to two Baltic Sea cables disrupted internet services between Finland, Germany, and Lithuania, with suspicions of foreign interference [2]. In December 2024 the EstLink2 power cable between Finland and Estonia was severed, prompting an investigation [3]. In February 2025, the C-Lion1 cable between Finland and Germany suffered damage near Gotland Island [1]. The European Commission recently addressed the need for stronger monitoring and security measures in its Joint Communication on submarine cable resilience [4].
@@
Detecting and localizing vessels near submarine cables is thus crucial to prevent disruptions in international data communications. Existing vessel localization methods use various sensing techniques, each with strengths and limitations. Yu et al. [5] demonstrated vessel detection with synthetic aperture radar (SAR), which enables wide-area surveillance in all weather conditions. However, SAR struggles with real-time processing and false alarms from sea clutter. Wawrzyniak et al. [6] used video surveillance in ports, applying vessel detection algorithms to track ships, but these systems are limited by lighting conditions and range. Similarly, Xie et al. [7] improved ship detection in multispectral satellite imagery with machine learning (ML), enhancing performance under mist and clouds, though severe weather and nighttime conditions significantly reduce accuracy.
@@
Distributed Acoustic Sensing (DAS) overcomes many limitations of traditional methods by repurposing submarine telecommunication cables as virtual sensor arrays [8]. Unlike SAR and optical imagery, DAS enables continuous real-time monitoring with high spatial and temporal resolution, independently of lighting and weather conditions [9]. It also operates without cooperative systems like the Automatic Identification System (AIS), allowing the detection of “dark” ships involved in unauthorized activities [10]. By making use of existing infrastructure, DAS provides wide-area coverage at lower costs without requiring dedicated hardware like hydrophones [10, 8]. DAS detects ships by analyzing their acoustic signatures, such as broadband noise from machinery and tonal signals from propellers. These characteristics have been studied for vessel classification, speed estimation, and maritime traffic monitoring [11]. However, existing methods must be adapted to the specific signal-processing requirements of DAS. Despite advancements, significant gaps remain in the DAS literature for vessel detection and localization, specially those oriented to submarine cable protection. Many studies use limited datasets, lack diversity in vessel types and environmental conditions, or focus on proof-of-concept setups without real-world validation.
@@ II Previous Work
DAS data present significant challenges due to its high-dimensionality, noise levels, and large volume, often surpassing the capabilities of traditional data storage and signal processing capabilities [12, 13, 14]. On the other hand, advanced techniques for data interpretation such as artificial intelligence and ML methods have demonstrated potential in overcoming these challenges for vessel detection [10, 14], seismic monitoring [15], and acoustic recognition [16, 17]. Malaprade et al.[10] illustrated the feasibility of ML-based ship detection using frequency band energy plots but were limited by a small dataset (only seven crossings) and sparse experimental details. Thiem et al.[9] localized a single passenger ship in a Norwegian fjord using noise-reduction filtering and traveltime inversion, showing good agreement with AIS but focusing on just one vessel type and not assessing other factors like tides or weather. Rivet et al.[14] estimated tanker trajectories off Toulon, France, leveraging Doppler effects and simulation models; while results aligned with observations, the study considered only one vessel and lacked broader environmental variability. Chen et al.[18] deployed a photonic integrated sensing and communication system in the Pearl River estuary, effectively tracking a ferry route, though only one vessel type and route were examined. Wienecke et al.[19] detected bottom trawlers via seabed vibrations up to 2.5km2.5\,km from the cable, but tested no multi-vessel scenarios. Similarly, Drylerakis et al.[16] used PCA-based denoising to improve signals from a single passing vessel.
@@
Other works emphasize shorter cable deployments or more constrained tests. Dias et al.[20] used a 500m500\,m cable in shallow waters to detect artificial ship noise, while Douglass et al.[21] captured broadband acoustic signals along a 3.5km3.5\,km long cable and compared DAS responses to hydrophones. Paap et al.[22] demonstrated automatic detection and localization for three specific vessels, achieving good performance but not exploring more diverse vessel types. Shao et al.[23] used Doppler shifts to track ships in a short 180m180\,m river cable deployment; the method aligned well with GPS but only covered three vessel passages.
@@
Very recently, Huang et al.[24] proposed DASHip, a large-scale labeled dataset for ship detection based on AIS-guided annotation of 55,875 events. They reported strong detection performance (73.6% joint detection ratio) using an object visual detection algorithm (YOLO), yet important higher-frequency acoustic cues (10200Hz10--200\,Hz) were sacrificed by downsampling to 10Hz10\,Hz, and their wake-centric approach could not identify vessels before cable crossings. This domain gap between clean labels and real-world conditions potentially reduces generalization[25, 8].
@@
Beyond direct ship detection, Landrø et al.[8] analyzed various acoustic phenomena, including seismic events and whale calls, using an Arctic cable. They tracked an 86-meter cargo ship via hyperbolic wavefront analysis and near-field beamforming. Stork et al.[15] likewise demonstrated microseismic and whale vocalization detection on a 2km2\,km cable, highlighting the versatility of DAS for multi-purpose acoustic monitoring. While such studies do not center on vessel detection, they confirm DAS’s broad applicability in ocean acoustic sensing.
@@ III Distributed Acoustic Sensing in Submarine Cables
Optical fiber-based distributed acoustic sensors are advanced sensing systems that use optical fibers as the primary medium for detecting and monitoring dynamic acoustic signals along their entire length. Unlike traditional sensor networks, DAS systems convert a conventional optical fiber into a vast array of virtual sensors that can measure dynamic strain changes along its entire length [26]. Conventional DAS technology probes the fiber-optic cable with a coherent laser pulse, and measures changes in the phase of the returning Rayleigh back-scatter signal time-series. Optical phase shifts in the back-scattered signal are proportional to longitudinal deformation in the fiber and can be translated into local strain changes across a fiber segment (termed gauge length $L$) by integration. The primary issue with the conventional phase-measuring approach is that a substantial fraction of positions along the fiber exhibit zero or near-zero intensity. These low-intensity points (typically called “fading points”) in the back-scattered signal lead to a defective determination of the phase of the local back-scattered field, implying inhomogeneous sensitivity variations across the DAS array [27]. Phase-measuring DAS systems generally available in the market achieve a fully distributed measurement of vibrations of up to several kHz, with a resolution of a few meters and measurement ranges of a few tens of kilometers. Optical coding methods are one of the preferred approaches for extending the measurement range of DAS systems. Among these, optical pulse compression reflectometry [28] has been widely developed, allowing measurement ranges well exceeding 100km100\,km [29].
Regardless of the interrogation method, DAS sensitivity depends on the optical fiber type, installation conditions, and local fiber-ground coupling. In existing installations, these factors are essentially not controllable; thus, effective data processing algorithms must account for them. Recent advances in signal processing and machine learning techniques have enabled improved handling of coupling variations, significantly enhancing DAS performance and making it a powerful tool for complex, large-scale monitoring tasks [30].
One of the mainstream applications of DAS is the monitoring of submarine cables for various purposes. In particular, DAS technology has proven effective in the surveillance of both communication and energy transmission submarine cables [31, 10]. Traditional offshore monitoring methods typically rely on fixed platforms, buoys, or ocean-bottom point sensors, offering limited spatial coverage. This constraint hinders the detection of dynamic and spatially diverse events, such as maritime activity involving vessels, large marine mammals, or other wildlife, occurring at the sea surface or seabed.
@@ IV-A Geographical and Instrumentation Details
The dataset was recorded using a 28km28\,km long pre-existing ocean-bottom fiber-optic cable in the Southern Bight of the North Sea offshore Zeebrugge, Belgium (see Fig. 1). The fiber-optic cable was originally installed to monitor a power cable from an offshore wind farm facility. The cable is buried between 1.4m1.4\,m and 7.2m7.2\,m below the seafloor, with an average burial depth of 4.7m4.7\,m. The Alcatel OptoDAS interrogator [32] was used for strain data acquisition. It is a phase-measuring DAS system that uses optical pulse compression reflectometry to extend the measurement range. The gauge length in this case is $L\approx 10\,m$, which is significantly larger than the nominal optical resolution of the system (1m1\,m), related to the bandwidth of the optical probing signal used. This significantly reduces the sensitivity issues related to fading points.
@@ IV-B1 AIS Data
In this study, AIS data were employed primarily for geographically labeling the DAS data, providing ground-truth vessel positions, which are essential for both training and evaluation purposes [33]. The raw AIS data, provided by the cable owner, included Maritime Mobile Service Identity (MMSI) numbers, UTC timestamps, vessel positions (longitude and latitude), course, speed, and heading for 745 unique vessels, totaling 64,417 reported positions. We further augmented this dataset by web-scraping static vessel information (name, IMO number, type, dimensions, tonnage), although dynamic details like navigational status and draught were unavailable. Analysis of vessel types revealed 45 distinct categories, with cargo ships being the most prevalent in the monitored area, alongside notable presences of fishing vessels and tugs (see Table I for detailed statistics).
@@
A significant challenge encountered when utilizing AIS data for machine learning applications, including ours, is the typically low and variable update rate [34, 33]. While ITU recommendations suggest reporting intervals under 30 seconds for most moving vessels [35], our dataset exhibited much lower frequencies, with $88\%$ of vessels reporting only every 1 to 3 minutes (see Fig. 2 for details). Considering the average vessel speed of 10.2 knots observed in the dataset (Fig. 3), such infrequent updates can lead to significant positional uncertainty (approximately 315 meters per minute in average), preventing accurate labeling. To address this limitation, we applied linear interpolation between reported positions with a 1-second resolution, excluding intervals with vessels not reporting during 60 minutes, and incorporating additional heuristic rules to minimize the generation of erroneous trajectories. This interpolation process resulted in a significantly denser dataset of 1,207,446 position entries used for subsequent analysis.
@@ IV-B2 Geometrical + bathymetry data acquisition
Accurate knowledge of the fiber-optic cable’s geographical position and the surrounding bathymetry is critical for DAS-based location-dependent monitoring. Variations in seafloor morphology and cable burial conditions directly impact the coupling between the cable and the seabed, significantly influencing the quality and characteristics of the recorded acoustic signals [36, 37]. Furthermore, precise fiber localization is essential for accurately positioning detected events, such as vessel passages. While cable deployment aims to follow a specific trajectory, the final resting position on the seafloor can deviate, and its burial status may change over time [36].
For this study, we utilized precise geometrical and bathymetry measurements for the full 28km28\,km cable length, provided directly by the cable owner. This dataset offers significantly higher precision than publicly available sources like EMODNet [38] (which has a resolution of 115 meters). A comparison, shown in Fig. 4, reveals depth discrepancies of up to 6 meters between the owner’s data and EMODNet along the cable path used in this work.
@@ IV-D Data Availability
Due to data-owner restrictions, neither the original raw differential-strain recordings nor precise geographical localization can be shared publicly. Instead, we have released the Marlinks-NS DAS dataset in Zenodo [39], that includes the complete set of processed data for the experiments in this work. It comprises the feature vectors and their corresponding ground-truth labels (timestamp, closest-vessel distance information, and vessel related metadata) in a structured hdf5 archive. To ease data distribution and reproducibility, we have also generated a GitHub repository (available at https://github.com/UAH-PSI/das-vessel-detection) that provides the data description, initial source code for loading and processing the data, as well as a small showcase data subset.
@@ V-A DAS Signal Preprocessing
  • Data scaling: Adjusts the differential phase measurements based on system gain and calibration coefficients, ensuring subsequent steps operate on signals with consistent amplitude levels.
@@
  • Spike detection and removal: Identifies and corrects sudden anomalies (or “spikes”) that can arise from large, rapid signal changes or noise bursts. These spikes are replaced or smoothed to prevent phase unwrapping errors.
  • Time integration: Integrates the unwrapped phase over time to recover the cumulative strain in the fiber. This step yields a direct strain signal rather than a differential measurement.

### v3

@@ I Introduction
Submarine cables, essential for global internet connectivity, energy transmission, and communication, are particularly vulnerable to accidental damage (e.g., fishing nets, ship anchors) and sabotage linked to geopolitical tensions [euromaidan_baltic_2025]. Recent incidents highlight this risk: in November 2024, damage to two Baltic Sea cables disrupted internet services between Finland, Germany, and Lithuania, with suspicions of foreign interference [cbsnews_baltic_cables_2024]. In December 2024 the EstLink2 power cable between Finland and Estonia was severed, prompting an investigation [navalnews_estlink2_2024]. In February 2025, the C-Lion1 cable between Finland and Germany suffered damage near Gotland Island [euromaidan_baltic_2025]. The European Commission recently addressed the need for stronger monitoring and security measures in its Joint Communication on submarine cable resilience [eu_submarine_cables2025].
@@
Detecting and localizing vessels near submarine cables is thus crucial to prevent disruptions in international data communications. Existing vessel localization methods use various sensing techniques, each with strengths and limitations. Yu et al. [Yu_2021] demonstrated vessel detection with synthetic aperture radar (SAR), which enables wide-area surveillance in all weather conditions. However, SAR struggles with real-time processing and false alarms from sea clutter. Wawrzyniak et al. [Wawrzyniak_2019] used video surveillance in ports, applying vessel detection algorithms to track ships, but these systems are limited by lighting conditions and range. Similarly, Xie et al. [Xie_2020] improved ship detection in multispectral satellite imagery with machine learning (ML), enhancing performance under mist and clouds, though severe weather and nighttime conditions significantly reduce accuracy.
@@
Distributed Acoustic Sensing (DAS) has emerged as a transformative technology enabling dense spatial sampling of acoustic signals along fiber-optic cables, and has seen increasing applications across geophysics, infrastructure monitoring, and environmental sensing. Recent reviews such as [cheng2024photonic] provide an overview of the evolution, challenges, and potentials of DAS across different domains. In the paper context, DAS overcomes many limitations of traditional methods by repurposing submarine telecommunication cables as virtual sensor arrays [Landro_2022]. Unlike SAR and optical imagery, DAS enables continuous real-time monitoring with high spatial and temporal resolution, independently of lighting and weather conditions [Thiem_2023]. It also operates without cooperative systems like the Automatic Identification System (AIS), allowing the detection of “dark” ships involved in unauthorized activities [Malaprade2019]. By making use of existing infrastructure, DAS provides wide-area coverage at lower costs without requiring dedicated hardware like hydrophones [Malaprade2019, Landro_2022]. DAS detects ships by analyzing their acoustic signatures, such as broadband noise from machinery and tonal signals from propellers. These characteristics have been studied for vessel classification, speed estimation, and maritime traffic monitoring [Malinowski_2002]. However, existing methods must be adapted to the specific signal-processing requirements of DAS. Despite advancements, significant gaps remain in the DAS literature for vessel detection and localization, specially those oriented to submarine cable protection. Many studies use limited datasets, lack diversity in vessel types and environmental conditions, or focus on proof-of-concept setups without real-world validation.
@@
To the best of our knowledge, this is the first study to systematically address vessel detection and distance estimation using machine learning in submarine cable DAS data, specifically focused in a submarine cable protection task, at a much larger scale than any other previous work.
@@ II Previous Work
DAS data present significant challenges due to its high-dimensionality, noise levels, and large volume, often surpassing the capabilities of traditional data storage and signal processing capabilities [Nur_DAS_storage2024, dong2020dassa, Rivet_2021]. On the other hand, advanced techniques for data interpretation such as artificial intelligence and ML methods have demonstrated potential in overcoming these challenges for vessel detection [Malaprade2019, Rivet_2021], seismic monitoring [Stork_2020], and acoustic recognition [Drylerakis2024source, zhan2024application]. Malaprade et al.[Malaprade2019] illustrated the feasibility of ML-based ship detection using frequency band energy plots but were limited by a small dataset (only seven crossings) and sparse experimental details. Thiem et al.[Thiem_2023] localized a single passenger ship in a Norwegian fjord using noise-reduction filtering and traveltime inversion, showing good agreement with AIS but focusing on just one vessel type and not assessing other factors like tides or weather. Rivet et al.[Rivet_2021] estimated tanker trajectories off Toulon, France, leveraging Doppler effects and simulation models; while results aligned with observations, the study considered only one vessel and lacked broader environmental variability. Chen et al.[Chen_2022] deployed a photonic integrated sensing and communication system in the Pearl River estuary, effectively tracking a ferry route, though only one vessel type and route were examined. Wienecke et al.[wienecke2023new] detected bottom trawlers via seabed vibrations up to 2.5km2.5\,km from the cable, but tested no multi-vessel scenarios. Similarly, Drylerakis et al.[Drylerakis2024source] used PCA-based denoising to improve signals from a single passing vessel.
@@
Other works emphasize shorter cable deployments or more constrained tests. Dias et al.[dias2014] used a 500m500\,m cable in shallow waters to detect artificial ship noise, while Douglass et al.[Douglass_2023] captured broadband acoustic signals along a 3.5km3.5\,km long cable and compared DAS responses to hydrophones. Paap et al.[paap2025leveragin] demonstrated automatic detection and localization for three specific vessels, achieving good performance but not exploring more diverse vessel types. Shao et al.[shao2025tracking] used Doppler shifts to track ships in a short 180m180\,m river cable deployment; the method aligned well with GPS but only covered three vessel passages.
@@
Very recently, Huang et al.[Huang_2025] proposed DASHip, a large-scale labeled dataset for ship detection based on AIS-guided annotation of 55,875 events. They reported strong detection performance (73.6% joint detection ratio) using an object visual detection algorithm (YOLO), yet important higher-frequency acoustic cues (10200Hz10--200\,Hz) were sacrificed by downsampling to 10Hz10\,Hz, and their wake-centric approach could not identify vessels before cable crossings. This domain gap between clean labels and real-world conditions potentially reduces generalization[Karasalo_2017, Landro_2022].
@@
Beyond direct ship detection, Landrø et al.[Landro_2022] analyzed various acoustic phenomena, including seismic events and whale calls, using an Arctic cable. They tracked an 86-meter cargo ship via hyperbolic wavefront analysis and near-field beamforming. Stork et al.[Stork_2020] likewise demonstrated microseismic and whale vocalization detection on a 2km2\,km cable, highlighting the versatility of DAS for multi-purpose acoustic monitoring. While such studies do not center on vessel detection, they confirm DAS’s broad applicability in ocean acoustic sensing.
@@ III Distributed Acoustic Sensing in Submarine Cables
Optical fiber-based distributed acoustic sensors are advanced sensing systems that use optical fibers as the primary medium for detecting and monitoring dynamic acoustic signals along their entire length. Unlike traditional sensor networks, DAS systems convert a conventional optical fiber into a vast array of virtual sensors that can measure dynamic strain changes along its entire length [Zhan_2020]. Conventional DAS technology probes the fiber-optic cable with a coherent laser pulse, and measures changes in the phase of the returning Rayleigh back-scatter signal time-series. Optical phase shifts in the back-scattered signal are proportional to longitudinal deformation in the fiber and can be translated into local strain changes across a fiber segment (termed gauge length $L$) by integration. The primary issue with the conventional phase-measuring approach is that a substantial fraction of positions along the fiber exhibit zero or near-zero intensity. These low-intensity points (typically called “fading points”) in the back-scattered signal lead to a defective determination of the phase of the local back-scattered field, implying inhomogeneous sensitivity variations across the DAS array [gabai2016sensitivity]. Phase-measuring DAS systems generally available in the market achieve a fully distributed measurement of vibrations of up to several kHz, with a resolution of a few meters and measurement ranges of a few tens of kilometers. Optical coding methods are one of the preferred approaches for extending the measurement range of DAS systems. Among these, optical pulse compression reflectometry [Zou_2015] has been widely developed, allowing measurement ranges well exceeding 100km100\,km [Waagaard_2021].
Regardless of the interrogation method, DAS sensitivity depends on the optical fiber type, installation conditions, and local fiber-ground coupling. In existing installations, these factors are essentially not controllable; thus, effective data processing algorithms must account for them. Recent advances in signal processing and machine learning techniques have enabled improved handling of coupling variations, significantly enhancing DAS performance and making it a powerful tool for complex, large-scale monitoring tasks [munoz2022enhancing].
One of the mainstream applications of DAS is the monitoring of submarine cables for various purposes. In particular, DAS technology has proven effective in the surveillance of both communication and energy transmission submarine cables [Brenne2019, Malaprade2019]. Traditional offshore monitoring methods typically rely on fixed platforms, buoys, or ocean-bottom point sensors, offering limited spatial coverage. This constraint hinders the detection of dynamic and spatially diverse events, such as maritime activity involving vessels, large marine mammals, or other wildlife, occurring at the sea surface or seabed.
@@ IV-A Geographical and Instrumentation Details
The dataset was recorded using a 28km28\,km long pre-existing ocean-bottom fiber-optic cable in the Southern Bight of the North Sea offshore Zeebrugge, Belgium (see Fig. 1). The fiber-optic cable was originally installed to monitor a power cable from an offshore wind farm facility. The cable is buried between 1.4m1.4\,m and 7.2m7.2\,m below the seafloor, with an average burial depth of 4.7m4.7\,m, which provides a good acoustic coupling between the seabed vibrations and the fiber [paap2025leveragin]. The Alcatel OptoDAS interrogator [OptoDAS] was used for strain data acquisition. It is a phase-measuring DAS system that uses optical pulse compression reflectometry to extend the measurement range. The gauge length in this case is $L\approx 10\,m$, which is significantly larger than the nominal optical resolution of the system (1m1\,m), related to the bandwidth of the optical probing signal used. This significantly reduces the sensitivity issues related to fading points.
@@ IV-B1 AIS Data
In this study, AIS data were employed primarily for geographically labeling the DAS data, providing ground-truth vessel positions, which are essential for both training and evaluation purposes [emmens2021promisesAIS]. The raw AIS data, provided by the cable owner, included Maritime Mobile Service Identity (MMSI) numbers, UTC timestamps, vessel positions (longitude and latitude), course, speed, and heading for 745 unique vessels, totaling 64,417 reported positions. We further augmented this dataset by web-scraping static vessel information (name, IMO number, type, dimensions, tonnage), although dynamic details like navigational status and draught were unavailable. Analysis of vessel types revealed 45 distinct categories, with cargo ships being the most prevalent in the monitored area, alongside notable presences of fishing vessels and tugs (see Table I for detailed statistics).
@@
A significant challenge encountered when utilizing AIS data for machine learning applications, including ours, is the typically low and variable update rate [Harati-Mokhtari_Wall_Brooks_Wang_2007AIS, emmens2021promisesAIS]. While ITU recommendations suggest reporting intervals under 30 seconds for most moving vessels [ITU_R_M1371_5AIS], our dataset exhibited much lower frequencies, with $88\%$ of vessels reporting only every 1 to 3 minutes (see Fig. 2 for details). Considering the average vessel speed of 10.2 knots observed in the dataset (Fig. 3), such infrequent updates can lead to significant positional uncertainty (approximately 315 meters per minute in average), preventing accurate labeling. To address this limitation, we applied linear interpolation between reported positions with a 1-second resolution, excluding intervals with vessels not reporting during 60 minutes, and incorporating additional heuristic rules to minimize the generation of erroneous trajectories. Given the update rate distribution and the good results achieved, we believe the linear interpolation is good enough for the explored scenario, although in more demanding situations, alternative interpolation methods should be adopted [guo2021improved] (more details in the supplementary material Web page). This interpolation process resulted in a significantly denser dataset of 1,207,446 position entries used for subsequent analysis.
@@ IV-B2 Geometrical + bathymetry data acquisition
Accurate knowledge of the fiber-optic cable’s geographical position and the surrounding bathymetry is critical for DAS-based location-dependent monitoring. Variations in seafloor morphology and cable burial conditions directly impact the coupling between the cable and the seabed, significantly influencing the quality and characteristics of the recorded acoustic signals [sladen2019distributed, Mata_Flores_2023]. Furthermore, precise fiber localization is essential for accurately positioning detected events, such as vessel passages. While cable deployment aims to follow a specific trajectory, the final resting position on the seafloor can deviate, and its burial status may change over time [sladen2019distributed].
For this study, we utilized precise geometrical and bathymetry measurements for the full 28km28\,km cable length, provided directly by the cable owner. This dataset offers significantly higher precision than publicly available sources like EMODNet [emodnet2024] (which has a resolution of 115 meters). A comparison, shown in Fig. 4, reveals depth discrepancies of up to 6 meters between the owner’s data and EMODNet along the cable path used in this work.
@@ IV-D Data Availability
Due to data-owner restrictions, neither the original raw differential-strain recordings nor precise geographical localization can be shared publicly. Instead, we have released the Marlinks-NS DAS dataset in Zenodo [ramirez2024dasvesseldataset], that includes the complete set of processed data for the experiments in this work. It comprises the feature vectors and their corresponding ground-truth labels (timestamp, closest-vessel distance information, and vessel related metadata) in a structured hdf5 archive. To ease data distribution and reproducibility, we have also generated a GitHub repository (available at https://github.com/UAH-PSI/das-vessel-detection) that provides the data description, initial source code for loading and processing the data, as well as a small showcase data subset.
@@ V-A DAS Signal Preprocessing
  • Data scaling: Adjusts the raw differential phase measurements based on system gain and calibration coefficients, ensuring subsequent steps operate on signals with consistent amplitude levels.
  • Spike detection and removal: Removes impulsive spikes using an adaptive thresholding method, where samples exceeding a sliding-window mean plus an offset are set to zero before integration.
@@
  • Time integration: Integrates the unwrapped phase over time and scale it to recover the cumulative strain in the fiber. This step yields a direct strain signal rather than a differential measurement. Fig. 8.a shows an example of a strain signal acquired during a vessel crossing above the cable.
Strain-rate and strain are linearly related in the frequency domain (essentially involving a multiplication of the strain-rate spectrum by $1/j\omega)$, so that the use of either magnitude would simply lead to changes in model tuning, and being related by a linear operation, it should definitely not affect the system performance.